1. through:(3,2), slope=2
y=m.x+b
2.through:(-2,1), slope=-3/2
3. through:(-1,2), slope=-7
4. through:(-2,2), slope=-5/2
5. through:(-5,-5), slope=2

Answer by Alan3354(69443)

1. through:(3,2), slope=2
y-y1 = m*(x-x1) where (x1,y1) is the point
---
y-2 = 2(x-3)
y = 2x - 4
===============
y=m.x+b
2.through:(-2,1), slope=-3/2
3. through:(-1,2), slope=-7
4. through:(-2,2), slope=-5/2
5. through:(-5,-5), slope=2
==================
Do 2 - 5 the same way.